# Plan Your Travel and Travel with Your Plan: Wide-Horizon Planning and Evaluation via LLM

### v1

@@ 4 Travel with Your Plan: Consistent Evaluation via Agent-based Simulation
Method Rule-based LLM-Judge Multi-Granularity Consistency
@@
4.1 Consistent Travel Experience Simulation
@@
The traveler follows the plan $p$ to begin the travel within the sandbox. We define the traveler’s state as $c_{n}=\{t,l,s,o,e\}$ at step $n$, where $t$, $l$, $s$, $o$, $e$ represent the current time, location, stamina, outlay, and the ongoing event. The traveler employs a policy $\pi(a_{n}|c_{1},c_{2},\cdots,c_{n-1},p)$, where action $a_{n}\in\mathcal{A}$ and action space $\mathcal{A}$ covers basic actions such as transiting, resting, dining, and sightseeing. Each action the traveler takes leads to a new event, which transits the previous state to the new one. Similar to ReAct yao2022react , the traveler agent first thinks over the current situation from the traveler’s perspective and then makes a decision on the next action.
@@
When coming to a new city, the traveler usually starts the travel at the train station or airport. The traveler can choose an action, in most cases, the first step is to proceed to the hotel. We utilize a map API to provide various modes of transportation as references. The traveler comprehensively considers stamina (being tired from the long flight), schedule (next event in the plan), and budget to make a decision. For example, if the traveler opts to take the metro to the hotel, the environment updates the traveler’s state with arrival time, new location (hotel), new stamina, etc. Among the events, the simulation of sightseeing events, due to the various experiences and interactions, presents a more significant challenge. Although the traveler agent can not physically visit the POI, we utilize a special sightseeing event agent to generate how the traveler would do the sightseeing by referring to the real experiences in the travel blog posts.
@@ 4.2 Multi-granularity Evaluation by Traveler
After the trip, the traveler undoubtedly has much to "share". We have the travelers give evaluations on the travel experience from five perspectives (see Appendix F.4 for more details), i.e., experience (ex), interest (it), arrangement (ar), stamina (st), and cost (co). We design a multi-granularity evaluation mechanism from three levels: the traveler reflects on the experience and rates the score after each POI visit, at the day’s conclusion, and after completing the entire journey.
@@
We synthesize a variety of traveler profiles, including 16 distinct types of travelers. Each differs in terms of group size, age, gender, stamina level, and preferences. We carefully select 7 Chinese cities that are ideal for tourism as destinations. We have 112 different distinct {traveler, destination, duration} combinations. See more details in Appendix F.3.
@@ Methods
For the baseline, we reuse the setup of preliminary experiments, including zero-shot long-horizon thinking and naive wide-horizon thinking. We add an additional baseline that we implement RL to train the models with Long/Artifact. setting using the same training dataset as the MAoP. This additional baseline is to compare the finetuned performance between the long-horizon thinking and the MAoP method. For the MAoP, we train the Qwen 2.5-7B yang2024qwen25 , Qwen 2.5-32B yang2024qwen25 , and Deepseek-R1-Distill Qwen-7B guo2025deepseekr1 as the strategist models and the planning models, respectively. We also include Deepseek-R1 guo2025deepseekr1 and Gemini-2.5-Pro-Exp-03-25 google2024gemini as additional zero-shot planning models. During the MAoP inference, the strategists only route no more than 8 aspects into the planning blueprint for fair comparison.
To simplify MAoP for one-step wide-horizon thinking, we distill the MAoT synthetic data to train Qwen 2.5-3B yang2024qwen25 , Llama 3.2-3B grattafiori2024llama3 , and Deepseek-R1-Distill Qwen-7B. The teachers in the distillation are the Deepseek-R1-Distill Qwen-7B as the strategist model and Gemini 2.5-Pro-Exp-0325 as the planning model. More training details can be checked in Appendix E.
Figure 3: The comprehensive comparison results between MAoP with the baseline methods.
@@
We evaluate the MAoP on our Travel-Sim to showcase how MAoP better deals with the $L^{3}$ planning problem via wide-horizon thinking. We introduce four metrics (with details in Appendix G): comprehensiveness (CPH), completeness (CPL), feasibility (FEA), and personalization (PER), where the last two metrics have been used in the preliminary experiment. The first two are rule-based metrics. Comprehensiveness (CPH) evaluates how much relevant information is effectively integrated from the long context into the final plan. Completeness (CPL) evaluates if the travel plan strictly follows the formatting instructions and basic constraints according to the artifactual criteria.
The last two metrics, feasibility (FEA) and personalization (PER), are based on the travel simulation. To evaluate the FEA, also named as Travel Plan Similarity Score (TPSS), we develop an algorithm that calculates the similarity between the trajectory from the travel plan and the trajectory from the simulated travel, as shown in Figure 5 and Algorithm 3. This metric measures the discrepancy between what is planned and the actual execution in the simulation. If the planned trajectory is similar to the simulated one, it means the plan is more feasible. The PER is associated with the feedback of the traveler agent after the simulated travel, indicating if the plan is personalized and suitable for this traveler. The PER is a comprehensive metric, including the evaluation from multiple dimensions and granularities: 1) Multi-dimension: The traveler evaluates the travel experience from five perspectives, i.e., experience (ex), interest (it), arrangement (ar), stamina (st), and cost (co). 2) Multi-granularity: the traveler conducts evaluation from three levels (feedback after visiting every POI, finishing the whole day, and finishing the whole journey). We aggregate the scores from various dimensions and granularities of travel experience, according to Equation 2, to calculate a final PER score.
@@ 5.2.1 Benchmark Performance
As illustrated in Figure 3, for the finetuned models in the first row, the MAoP demonstrates a substantial performance enhancement over the long-horizon thinking (RL w/ Long/Artifact.) baselines, achieving a remarkable 5% to 40% improvement across all planning models. Although trained with the same dataset, MAoP achieves higher scores than RL w/ Long/Artifact. in CPL by better constraint compliance and CPH by integrating more details from the long context to the plan. When compared with naive wide-horizon thinking (Zero-shot w/ Wide/Aritifact.), MAoP excels especially in stronger strategist models, as evidenced by higher FEA and PER scores, suggesting that the strategist plays an important role in the planning process with performance better than artifactual guidance. For advanced models as zero-shot planning models, e.g., Deepseek-R1 and Gemini-2.5, the strategist model also significantly boosts the performance in travel planning compared to zero-shot baselines.
5.2.2 Inference-Time Scaling Capability of the Strategist Model
Figure 4: We experiment two strategist models (Qwen 7B & R1-Distill 7B) and two planning models (Qwen 7B & Gemini) to showcase the scaling capability of the strategist models.
As shown in Figure 4, we find that strategist models can consistently improve the performance by scaling up more considered aspects. We find that even if the model size is the same, the stronger strategist with thinking (R1-Distill 7B) has better scalability, especially for advanced planning models like Gemini 2.5. In contrast, the Qwen 2.5-7B model has limited scalability, because it cannot effectively route a suitable planning blueprint when dealing with increasing aspects.
@@ 5.2.3 MAoP Distillation
As shown in Table 3, distilling R1-Distill 7B (s.) + Gemini 2.5-Pro-Exp-0325 (p.) enables 3B sized models to outperform the MAoP combinations even with larger model sizes. For R1-Distill 7B with thinking mode, we also distill the thinking part of planning from the Gemini. Compared to the MAoP combination of the same models, R1-Distill 7B (Distill) achieves better performance. These results demonstrate that distillation from advanced models to smaller models achieves substantial performance improvements, with the enhancement becoming increasingly pronounced as the capability gap between teacher and student models widens, even surpassing the MAoP training models.
5.2.4 Emergence of Spontaneous Behaviors in Consistent Travel Simulation
Instead of strictly obeying the plan, the travel agent refers to the plan and dynamically decides their own itinerary based on various other factors, e.g., stamina, environmental changes, etc. Similar to real-world scenarios, the itinerary is consistent, and subsequent activities are influenced by previous events. Therefore, we observe the emergent behaviors that the traveler agent spontaneously adjusts the predetermined plan, which is not pre-programmed. As the trajectories shown in Figure 5, this elderly couple chooses not to have dinner in Nanmen Lamb Hot Pot because they are too tired to travel to another place, especially since they just suffered from a long train trip to Beijing this morning. In real travel scenarios, when one event goes awry—such as running out of time, excessive fatigue, or attractions being closed—it disrupts the plans for all subsequent activities. This indicates that neglecting consistency issues in travel plan evaluations can lead to significant deviations from real-world scenarios. Travel-Sim surpasses other evaluation benchmarks in assessing consistency by conducting simulations within a real-world sandbox environment.
Figure 5: The example of the trajectories from the generated plan and the trajectory from the simulation. The trajectory of the simulation is not always consistent with the planned one because the traveler agent can change the subsequent itinerary based on the current situation. The FEA score is to calculate the similarity of these two trajectories.
@@ 6 Related Work
The planning capabilities of LLMs have become a key research focus due to their potential as autonomous problem-solving agents wei2023chainofthoughtpromptingelicitsreasoning ; yao2022react ; Wang_2024 . Recent studies have advanced these capabilities through developments in task decomposition, multi-step reasoning yao2023treethoughtsdeliberateproblem ; Besta_2024 , and adaptive planning sun2023adaplanneradaptiveplanningfeedback ; jutrasdubé2024adaptiveplanninggenerativemodels . While these planning algorithms have shown promising results hao2023reasoninglanguagemodelplanning ; hu2024treeplannerefficientcloselooptask , their planning scenarios are limited to simple tasks with a single objective function. For complex real-world planning like travel planning, previous researchers focus on how to preprocess the various information and constraints for LLMs easier to understand but few of them explore how to enhance the planning ability in this complex scenario.
@@ 7 Conclusion
We propose MAoP to enhance wide-horizon thinking for solving $L^{3}$ planning problems, thus addressing the complex challenge of travel planning. In addition, we propose Travel-Sim, an evaluation benchmark that leverages agent-based simulation to offer consistent and multi-granularity evaluation in travel planning scenarios. Our contributions advance the wide-horizon thinking capabilities of LLMs in real-world planning and offer novel insights for evaluating sophisticated scenarios through agent-based simulation.
@@ A.1.1 MAoP Limitations
Though MAoP significantly enhances wide-horizon thinking with deliberate pre-planning by the strategist, MAoP still exhibits a few limitations: 1) MAoP requires the collaboration between the strategist model and planning model, and the conduction of aspect-aware thinking in multiple turns of dialogue. This complicated process costs more inference time than directly CoT. To alleviate this, we propose that the MAoP distillation can compress this complicated process, gaining similar efficiency compared to CoT. 2) We find that weak strategists, especially sizes smaller than 7B, can cause severe performance degradation due to low-quality planning blueprint and poor inference-time scalability. The recommended way to enhance wide-horizon thinking ability for smaller models is to distill the outputs of the advanced model after MAoP training.
@@ A.1.2 Travel-Sim Limitations
The Travel-Sim is the pioneer in evaluating with agent-based simulation, offering a novel solution for complex scenario evaluation, e.g., travel planning. Agent-based simulation and evaluation have several limitations that can be discussed: 1) Similar to other LLM-based evaluation benchmarks, the "traveler", which is an LLM, has the potential bias and unreproducibility over the evaluation. To mitigate this limitation, we stipulate the agent model to be Gemini-2.5-Pro-Exp-0325 [21] using greedy decoding (0 temperature) for reproducibility. We also verify the consistency between LLM-based and human-based evaluation in the Appendix F.5, showcasing the 92% consistency in PER score. 2) Although Travel-Sim takes the consistency and real-world information into account, we do not include all the emergencies that may happen in the real trip, especially some force majeure and human factors. For example, bad weather or delays caused by carelessness also significantly influence the trip. 3) Although Travel-Sim offers consistent and multi-granularity evaluation, the whole evaluation process is time-consuming and costs about 20K tokens per sample. To evaluate the 112 samples of the entire dataset, we cost around \$12 by consuming about 2M tokens in total.
@@ Appendix B License
By downloading or using the data, you understand, acknowledge, and agree to all the terms in the following agreement.
@@ Appendix C Preprocessing Framework
In this section, we introduce how to collect and preprocess the heterogeneous information as the "long context" for travel planning. This context information serves as the information source input both in the training dataset and Travel-Sim.
@@ C.1.2 Hotel Consultant Agent
The selection of the hotel is the key part of travel planning, which seldom previous researchers take into account. The hotel influences not only the accommodation experience but also the planning and logistics of the entire trip, concerning the hotel locations. We develop an LLM-based agent to serve as a hotel consultant agent to recommend the hotel based on travelers’ preferences.
@@
To be specific, in the inquiry phase, we consider four aspects, as shown in Table 5. After several proactive interactions, the agent will stop asking if it has collected enough user preferences. Besides the hotel preference, the hotel agent can leverage POI preferences and spatial information (POI clusters) from other modules. In the online searching phase, we use the search API provided by the Red Note (Xiaohongshu) app, a popular Chinese app focused on lifestyle content sharing. We use Red Note (Xiaohongshu) search API to search for blog posts related to hotel recommendations. In the map searching phase, we use the Amap (Gaode) Map API, a Chinese map app similar to Google Maps, to search for hotels near the most-visited locations. The probable most-visited locations are predicted by the agent based on the given information. In the conclusion phase, the agent will recommend the hotels provided by the blog posts and the map API by comprehensively taking the locations and preferences into account.

### v2

@@ 4 Causal Evaluation via Agent-based Simulation
Method Rule-based LLM-Judge Multi-Granularity Causality
@@
4.1 Travel Experience Simulation
@@
The traveler follows the plan $p$ to travel in the sandbox. We define the traveler’s state as $c_{n}=\{t,l,s,o,e\}$ at step $n$, where $t$, $l$, $s$, $o$, $e$ represent the current time, location, stamina, outlay, and the ongoing event. The traveler employs a policy $\pi(a_{n}|c_{1},c_{2},\cdots,c_{n-1},p)$, where action $a_{n}\in\mathcal{A}$ and action space $\mathcal{A}$ covers basic actions such as transiting, resting, dining, and sightseeing. Each action the traveler takes leads to a new event, which transitions the previous state to the new one. Similar to ReAct yao2022react , the traveler agent first thinks over the current situation from the traveler’s perspective and then makes a decision on the next action.
@@
When coming to a new city, the traveler usually starts the travel at the train station or airport. The traveler can choose an action; in most cases, the first step is to proceed to the hotel. We utilize a map API to provide various modes of transportation as references. The traveler comprehensively considers stamina (being tired from the long flight), schedule (next event in the plan), and budget to make a decision. For example, if the traveler opts to take the metro to the hotel, the environment updates the traveler’s state with arrival time, new location (hotel), new stamina, etc. Among the events, the simulation of activities like sightseeing or shopping, due to the various experiences and interactions, presents a more significant challenge. Although the traveler agent can not physically visit the POI, we utilize a special event agent to generate how the traveler would do the sightseeing by referring to the real experiences in the travel blog posts.
@@ 4.2 Multi-granularity Evaluation by Traveler
To capture detailed feedback, we implement a multi-granularity evaluation mechanism. Travelers assess their experience across five core dimensions: experience (ex), interest (it), arrangement (ar), stamina (st), and cost (co) (detailed in Appendix F.4). These evaluations are collected at three distinct levels of granularity: per-POI (after each POI visit), per-day (at the conclusion of each day), and per-trip (upon completion of the entire journey).
@@
We construct a variety of traveler profiles, including 16 distinct types of travelers. Each differs in terms of group size, age, gender, stamina level, and preferences. We carefully select 7 Chinese cities that are ideal for tourism as destinations. We have 112 different distinct {traveler, destination, duration} combinations. See more details in Appendix F.3.
@@ Methods
For the baseline, we reuse the setup of preliminary experiments, including zero-shot long-horizon thinking and naive wide-horizon thinking. We add an additional baseline that we implement RL to train the models with Long/Artifact. setting using the same training dataset as the MAoP. This additional baseline is to compare the finetuned performance between the long-horizon thinking and the MAoP method. For the MAoP, we train the Qwen 2.5-7B yang2024qwen25 , Qwen 2.5-32B yang2024qwen25 , and Deepseek-R1-Distill Qwen-7B guo2025deepseekr1 as the strategists and the planners, respectively. We also include Deepseek-R1 guo2025deepseekr1 and Gemini-2.5-Pro-Exp-03-25 google2024gemini as additional zero-shot planners. During the MAoP inference, the strategists only route no more than 8 aspects into the planning blueprint for fair comparison.
To simplify MAoP for one-step wide-horizon thinking, we distill the MAoP synthetic data to train Qwen 2.5-3B yang2024qwen25 , Llama 3.2-3B grattafiori2024llama3 , and Deepseek-R1-Distill Qwen-7B. The teachers in the distillation are the Deepseek-R1-Distill Qwen-7B as the strategist and Gemini 2.5-Pro-Exp-0325 as the planner. More training details can be checked in Appendix E.
Figure 3: The comprehensive comparison results between MAoP and the baseline methods.
@@
We evaluate the MAoP on our Travel-Sim to showcase how MAoP better deals with the travel planning problem via wide-horizon thinking. We introduce four metrics (with details in Appendix G): comprehensiveness (CPH), completeness (CPL), feasibility (FEA), and personalization (PER), where the last two metrics have been used in the preliminary experiment. The first two are rule-based metrics. Comprehensiveness (CPH) evaluates how much relevant information is effectively integrated from the long context into the final plan. Completeness (CPL) evaluates if the travel plan strictly follows the formatting instructions and basic constraints according to the artifactual criteria.
The last two metrics, feasibility (FEA) and personalization (PER), are based on the travel simulation. To evaluate the FEA, also named Travel Plan Similarity Score (TPSS), we develop an algorithm that calculates the similarity between the trajectory from the travel plan and the trajectory from the simulated travel, as shown in Figure 5 and Algorithm 3. This metric measures the discrepancy between what is planned and the actual execution in the simulation. If the planned trajectory is similar to the simulated one, it means the plan is more feasible. The PER is associated with the feedback of the traveler agent after the simulated travel, indicating if the plan is personalized and suitable for this traveler. The PER is a comprehensive metric, including the evaluation from multiple dimensions and granularities: 1) Multi-dimension: The traveler evaluates the travel experience from five perspectives, i.e., experience (ex), interest (it), arrangement (ar), stamina (st), and cost (co). 2) Multi-granularity: the traveler conducts evaluation from three levels (feedback after visiting every POI, finishing the whole day, and finishing the whole journey). We aggregate the scores from various dimensions and granularities of travel experience, according to Equation 3, to calculate a final PER score.
@@ 5.2.1 Benchmark Performance
Figure 4: We experiment two strategists (Qwen 7B & R1-Distill Qwen-7B) and two planners (Qwen 7B & Gemini 2.5) to showcase the scaling capability of the strategists.
As illustrated in Figure 3, for the finetuned models in the first row, the MAoP demonstrates a substantial performance enhancement over the long-horizon thinking (RL w/ Long/Artifact.) baselines, achieving a remarkable 5% to 40% improvement across all planners. Although trained with the same dataset, MAoP achieves higher scores than RL w/ Long/Artifact. in CPL by better constraint compliance and CPH by integrating more details from the long context to the plan. When compared with naive wide-horizon thinking (Zero-shot w/ Wide/Aritifact.), MAoP excels especially in stronger strategists, as evidenced by higher FEA and PER scores, suggesting that the strategist plays an important role in the planning process with performance better than artifactual guidance. For advanced models as zero-shot planners, e.g., Deepseek-R1 and Gemini-2.5, the strategist also significantly boosts the performance in travel planning compared to zero-shot baselines.
5.2.2 Inference-Time Scaling Capability of the Strategist
As shown in Figure 4, we find that strategists can consistently improve the performance by scaling up more considered aspects. We find that even if the model size is the same, the stronger strategist with thinking (R1-Distill 7B) has better scalability, especially for advanced planners like Gemini 2.5. In contrast, the Qwen 2.5-7B model has limited scalability, because it cannot effectively route a suitable planning blueprint when dealing with increasing aspects.
@@ 5.2.3 MAoP Distillation
As shown in Table 3, distilling R1-Distill 7B (s.) + Gemini 2.5-Pro-Exp-0325 (p.) enables 3B sized models to outperform the MAoP combinations even with larger model sizes. For R1-Distill 7B with the thinking mode, we also distill the thinking part of planning from the Gemini. Compared to the MAoP combination of the same models, R1-Distill 7B (Distill) achieves better performance with distilled one-step wide-horizon thinking. These results demonstrate that distillation from advanced models to smaller models achieves substantial performance improvements even without the original multi-turn MAoP process. The enhancement becomes increasingly pronounced as the capability gap between teacher and student models widens, even surpassing the MAoP training models.
5.2.4 Emergence of Spontaneous Behaviors in Causal Travel Simulation
Given the causal nature of an itinerary, the traveler agent exhibits emergent behavior by spontaneously adjusting its predetermined plan in response to unfolding events. From trajectories in Figure 5, this elderly couple chooses not to have dinner in Nanmen Lamb Hot Pot because they are too tired to travel to another place, especially since they just suffered from a long train trip to Beijing this morning. This indicates that neglecting causal dependencies in real-world planning evaluation can lead to significant deviations from the real-world situation.
Figure 5: The simulation trajectory is not always consistent with the planned one because the traveler agent can change the subsequent itinerary based on the current situation. The FEA score is used to calculate the similarity of these two trajectories.
@@ 6 Related Work
The planning capabilities of LLMs have become a key research focus due to their potential as autonomous problem-solving agents wei2023chainofthoughtpromptingelicitsreasoning ; yao2022react ; Wang_2024 . Recent studies have advanced these capabilities through developments in task decomposition, multi-step reasoning yao2023treethoughtsdeliberateproblem ; Besta_2024 , and adaptive planning sun2023adaplanneradaptiveplanningfeedback ; jutrasdubé2024adaptiveplanninggenerativemodels . While these planning algorithms have shown promising results hao2023reasoninglanguagemodelplanning ; hu2024treeplannerefficientcloselooptask , their planning scenarios are limited to simple tasks with a single objective function. For complex real-world planning like travel planning, previous researchers focus on how to preprocess the various information and constraints to make LLM easier to understand, but few of them explore how to enhance the planning ability in this complex scenario.
@@ 7 Conclusion
We propose MAoP to enhance wide-horizon thinking for solving real-world planning problems with multifaceted constraints. In addition, we propose Travel-Sim, an evaluation benchmark that leverages agent-based simulation to offer causal and multi-granularity evaluation in travel planning scenarios. Our contributions advance the wide-horizon thinking capabilities of LLMs in real-world planning and offer novel insights for evaluating sophisticated scenarios through agent-based simulation.
@@ A.1.1 MAoP Limitations
Though MAoP significantly enhances wide-horizon thinking with deliberate pre-planning by the strategist, MAoP still exhibits a few limitations: 1) MAoP requires the collaboration between the strategist and planner, and the conduction of aspect-aware thinking in multiple turns of dialogue. This complicated process costs more inference time than directly using CoT. To alleviate this, we propose that the MAoP distillation can compress this complicated process, gaining similar efficiency compared to CoT. 2) We find that weak strategists, especially those with sizes smaller than 7B, can cause severe performance degradation due to low-quality planning blueprint and poor inference-time scalability. The recommended way to enhance wide-horizon thinking ability for smaller models is to distill the outputs of the advanced model after MAoP training.
@@ A.1.2 Travel-Sim Limitations
The Travel-Sim is the pioneer in evaluating with agent-based simulation, offering a novel solution for complex scenario evaluation, e.g., travel planning. Agent-based simulation and evaluation have several limitations that can be discussed: 1) Similar to other LLM-based evaluation benchmarks, the "traveler", which is an LLM, has the potential bias and unreproducibility over the evaluation. To mitigate this limitation, we stipulate the agent model to be Gemini-2.5-Pro-Exp-0325 [21] using greedy decoding (0 temperature) for reproducibility. We also verify the consistency between LLM-based and human-based evaluation in the Appendix F.5, showcasing the 92% consistency in PER score. 2) Although Travel-Sim takes causality and real-world information into account, we do not include all the emergencies that may happen in the real trip, especially some force majeure and human factors. For example, bad weather or delays caused by carelessness also significantly influence the trip. 3) Although Travel-Sim offers causal and multi-granularity evaluation, the whole evaluation process is time-consuming and costs about 20K tokens per sample. To evaluate the 112 samples of the entire dataset, we cost around \$12 by consuming about 2M tokens in total.
@@ Appendix B License
By downloading or using our open-source data, you understand, acknowledge, and agree to all the terms in the following agreement.
@@ Appendix C Preprocessing Framework
In this section, we introduce how to collect and preprocess the heterogeneous information as the "long context" for travel planning. This context information serves as the information source input for both the training dataset and Travel-Sim.
@@ C.1.2 Hotel Consultant Agent
The selection of the hotel is the key part of travel planning, which previous researchers seldom take into account. The hotel influences not only the accommodation experience but also the planning and logistics of the entire trip, concerning the hotel locations. We develop an LLM-based agent to serve as a hotel consultant agent to recommend hotels based on travelers’ preferences.
@@
To be specific, in the inquiry phase, we consider four aspects, as shown in Table 5. After several proactive interactions, the agent will stop asking if it has collected enough user preferences. Besides the hotel preference, the hotel agent can leverage POI preferences and spatial information (POI clusters) from other modules. In the online searching phase, we use the search API provided by the Red Note (Xiaohongshu) app, a popular Chinese app focused on lifestyle content sharing. We use the Red Note (Xiaohongshu) search API to search for blog posts related to hotel recommendations. In the map searching phase, we use the Amap (Gaode) Map API, a Chinese map app similar to Google Maps, to search for hotels near the most-visited locations. The probable most-visited locations are predicted by the agent based on the given information. In the conclusion phase, the agent will recommend the hotels provided by the blog posts and the map API by comprehensively taking the locations and preferences into account.